# Unusual switch from low-temperature T-quadratic resistivity in the underdoped pseudogap phase of cuprate superconductors to low-temperature T-linear resistivity in the overdoped strange-metal phase

### v1

@@ II.4 Electrical transport due to umklapp scattering from a spin excitation
It thus shows that the electrical resistance originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator.
@@ III Quantitative characteristics
Now we are ready to discuss the low-temperature electrical transport in the underdoped pseudogap phase. In Fig. 5, we plot the resistivity $\rho(T)$ in Eq. (45) as a function of temperature at doping $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line), where in the low-temperature region, a power-law resistivity appears over a wide doping range in the underdoped regime.
Figure 5: (Color online) Resistivity as a function of temperature at $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line). Inset: the numerical fit of the resistivity (black-dots) at $\delta=0.09$ with the fit form $\rho(T)=A_{2}T^{2}$.
@@
To explore this power-law behaviour of the low-temperature resistivity more clearly, the above results of $\rho(T)$ in Fig. 5 have been numerically fitted with the fit form $\rho(T)=A_{2}T^{2}$, and the fitted result of the low-temperature resistivity (black-dots) at $\delta=0.09$ is also plotted in Fig. 5 (inset), where the low-temperature resistivity is demonstrated clearly to be grown quadratically as the temperature is raised. Moreover, the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) $A_{2}$ grows as the doping is reduced. To see this doping dependence of the T-quadratic resistivity strength more clearly, we plot $A_{2}$ as a function of doping $\delta$ in Fig. 6a. For a better comparison, the corresponding transport experimental resultBarisic13 observed on the underdoped pseudogap phase of cuprate superconductors is also shown in Fig. 6b. It thus shows clearly that $A_{2}$ drops gradually as doping is enhanced in the underdoped regime, indicating that $A_{2}$ is roughly proportional to the inverse doping concentration. All these results are in qualitative agreement with the corresponding transport experimental results observed on the underdoped pseudogap phase of cuprate superconductors Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 . These findings in Fig. 5 and Fig. 6 in the underdoped pseudogap phase along with the recent studyMa23 on the low-temperature T-linear resistivity in the overdoped strange-metal phase therefore show that the electron umklapp scattering from a spin excitation responsible for the T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap. As a nature consequence, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
Figure 7: (Color online) Surface plot of the kernel function $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively.
@@
The low-temperature resistivity is mainly determined by the umklapp scattering rate $\gamma(\theta,T)$ in Eq. (40), however, this umklapp scattering rate $\gamma(\theta,T)$ is in turn directly connected with the kernel function $F(\theta,\theta^{\prime})$ (then the probability weight of the umklapp scattering) in Eq. (41). For the further understanding of the nature of the umklapp scattering between electrons by the exchange of the full effective spin propagator, we first analyse the exotic feature of the kernel function $F(\theta,\theta^{\prime})$. In Fig. 7, we plot $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where the probability weight of the electron umklapp scattering is very strongly anisotropic in momentum space. In particular, as in the case of the overdoped strange-metal phase Ma23 , the almost all of the probability weight of the electron umklapp scattering in the underdoped pseudogap phase is also concentrated at around the antinodal region, leading to that the strongest umklapp scattering occurs at around this antinodal region.
@@
Figure 8: Umklapp scattering rate $\gamma(T)$ at the antinode as a function of temperature for $\delta=0.09$.
@@
However, the weakest umklapp scattering appears at around the tips of the Fermi arcs, since the probability weight of the electron umklapp scattering there is almost none. Moreover, the strength of the electron umklapp scattering at around the nodal region is much weaker than that at around the antinodal region, since a very small amount of the probability weight of the electron umklapp scattering is concentrated at around the nodal region. This special momentum dependent distribution of the probability weight of the electron umklapp scattering in Fig. 7 together with the recent discussions of the momentum-dependent distribution of the probability weight of the electron umklapp scattering in the overdoped strange-metal phaseMa23 therefore confirm that in the whole doping regime, the dominant contribution to the scattering rate mainly arises from the umklapp scattering between the relatively slow electrons at around the antinodal region.
@@
We now turn to explore the unusual evolution of the umklapp scattering rate (40) with temperature in the underdoped pseudogap phase. We have made a series of calculations for the umklapp scattering rate $\gamma(T)$ at different Fermi angles, and the result of $\gamma(T)$ as a function of temperature for $\delta=0.09$ at the antinode is plotted in Fig. 8, where the numerical fit demonstrates that $\gamma(T)$ is purely T-quadratic in the low-temperature region, i.e., it increases quadratically with temperature as the temperature is increased. More specifically, although the magnitude of $\gamma(\theta,T)$ is strongly anisotropic in momentum space, the low-temperature T-quadratic behaviour of $\gamma(\theta,T)$ appears at an any given Fermi angle $\theta$. Comparing this result in Fig. 8 with the corresponding results of the low-temperature T-quadratic resistivity in Fig. 5, it thus shows that the low-temperature T-quadratic behaviour of $\gamma(T)$ together with the temperature region are the same as the corresponding behaviour and region in the resistivity $\rho(T)$, which therefore confirms that the low-temperature T-quadratic resistivity with the corresponding temperature region is mainly governed by the low-temperature T-quadratic umklapp scattering rate with the corresponding temperature region.
@@
Now we give an explanation to show why the low-temperature resistivity exhibits a dramatic switch from the T-linear resistivity in the overdoped strange-metal phase to the T-quadratic resistivity in the underdoped pseudogap phase? As the mentioned above, the nature of the umklapp scattering rate in Eq. (40) is mainly determined by the nature of the kernel function $F(\theta,\theta^{\prime})$ in Eq. (41). However, this kernel function $F(\theta,\theta^{\prime})$ in the underdoped pseudogap phase is proportional to the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (32). From the full spin propagator in Eq. (17), the full spin spectral function is derived straightforwardly as $A_{\rm spin}({\bf k},\omega)=A^{(1)}_{\rm spin}({\bf k},\omega)+A^{(2)}_{\rm spin}({\bf k},\omega)$, where the components $A^{(1)}_{\rm spin}({\bf k},\omega)$ and $A^{(2)}_{\rm spin}({\bf k},\omega)$ are given by,
@@
respectively. However, during the calculation, we have found that the spectral weight in the component $A^{(1)}_{\rm spin}({\bf k},\omega)$ is several orders of magnitude greater than the corresponding spectral weight in the component $A^{(2)}_{\rm spin}({\bf k},\omega)$, leading to that the electron umklapp scattering is mainly mediated by the spin excitations from the component of $A^{(1)}_{\rm spin}({\bf k},\omega)$. On the other hand, the spin excitations with the full spin excitation dispersion $\bar{\omega}_{1{\bf k}}$ around the AF wave vector ${\bf k}_{\rm A}=[\pm\pi,\pm\pi]$ have the largest density of states. With these special properties of the spin excitations in the underdoped pseudogap phase, the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (37) can be reduced approximately as,
@@
with $b_{\bf q}=d^{2}\bar{\omega}_{1{\bf q}}/(d^{2}q)$. The above results in Eq. (48) show that as in the case of the the overdoped strange-metal phaseMa23 , the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (47) scales with $p^{2}$, and then when the electron umklapp scattering kicks in, the energy scale is proportional to $\Delta^{2}_{p}$ due to the presence of this $p^{2}$ scaling in the full effective spin propagator (47). In this case, $T_{\rm scale}=\bar{b}\Delta^{2}_{p}$ can be identified as the low-temperature scale,
with the average value $\bar{b}=(1/N)\sum\limits_{{\bf q}\in\{{\bf k}_{\rm A}\}}b({\bf q})$ that is a constant at a given doping, where the summation ${\bf q}\in\{{\bf k}_{\rm A}\}$ is restricted to the extremely small area $\{{\bf k}_{\rm A}\}$ around the ${\bf k}_{\rm A}$ point of BZ. However, this low-temperature scale $T_{\rm scale}$ is strongly doping dependent. To see this doping dependence of $T_{\rm scale}$ more clearly, we plot $T_{\rm scale}$ as a function of doping in Fig. 9, where $T_{\rm scale}$ is relatively high at around the slightly underdoped region, then it decreases when the doping is increased in the heavily underdoped region, and is reduced to a very low temperature in the overdoped regime. Moreover, $T_{\rm scale}$ as a function of doping in Fig. 9 in the underdoped regime presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$ shown in Fig. 2b, suggesting a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase.
@@ IV Summary and discussion
Within the framework of the kinetic-energy-driven superconductivity, we have rederived the full spin propagator, the full charge-carrier propagator, and the full electron propagator in the normal-state of cuprate superconductors, where the spin self-energy (then the spin pseudogap) in the normal-state is obtained explicitly in terms of the collective charge-carrier mode in the particle-hole channel, and the charge-carrier self-energy (then the charge-carrier pseudogap) in the normal-state is obtained explicitly in terms of the spin excitation mode, while the electron self-energy (then the normal-state pseudogap) originated from the charge-carrier self-energy (then the charge-carrier pseudogap) is due to the charge-spin recombination. Moreover, we have also shown that (i) the spin excitation energy dispersion is anisotropically renormalized due to the momentum dependence of the spin pseudogap. In particular, the antinodal spin pseudogap effect is particularly notable in the slightly underdoped region, and then this antinodal spin pseudogap rapidly decreases with the increase of doping in the heavily underdoped region, eventually abrupt disappearing at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. By virtue of this full spin propagator and the reconstructed EFS, we have investigated the low-temperature electrical transport in the underdoped pseudogap phase of cuprate superconductors, where the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived within the framework of the Boltzmann transport theory. Our results show that the dominant contribution to the low-temperature electrical transport mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Moreover, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime, which suggests a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase. In the underdoped regime, the resistivity exhibts a T-quadratic behaviour in the low-temperature region below $T_{\rm scale}$, where the strength of the T-quadratic resistivity decreases with the increase of doping. However, in the overdoped regime, the resistivity is T-linear in the low-temperature region above $T_{\rm scale}$. The current findings together with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped strange-metal phase therefore show that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped strange-metal phase; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependence of the spin pseudogap lowers the density of stats of the spin excitations at around the antinodal region in the response to the intense umklapp scattering, which reduces the strength of the umklapp scattering from electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. As a natural consequence, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.

### v2

@@ II.4 Electrical transport due to umklapp scattering from a spin excitation
It thus shows that the electrical resistance originates from the umklapp scattering between electrons by the exchange of the full effective spin propagator.
@@ III Quantitative characteristics
Figure 5: (Color online) (a) Resistivity as a function of temperature at $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line). Inset: the numerical fit of the resistivity (black-dots) at $\delta=0.09$ with the fit form $\rho(T)=A_{2}T^{2}$. (b) The corresponding experimental result of the low-temperature T-quadratic resistivity for the underdoped HgBa2CuO4+δ taken from Ref. Mirzaei13, .
It should be emphasized that the above resistivity in Eq. (32) is obtained in the pure two-dimensional $t$-$J$ model (1) on a square lattice. However, for a clear comparison with the corresponding experimental data Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , the resistivity obtained in the above equation (32) should be renormalized by the distance between the adjacent copper-oxide layers as it has been done in Ref. Bergeron11, , with the interlayer lattice constant Bergeron11 that can be chosen as $d=0.5$nm. Now we are ready to discuss the low-temperature resistivity in the underdoped pseudogap phase. In Fig. 5a, we plot the renormalized resistivity $\rho(T)$ as a function of temperature at doping $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line), where in the low-temperature region, a power-law resistivity appears over a wide doping range in the underdoped regime. To explore this power-law behaviour of the low-temperature resistivity more clearly, the above results of $\rho(T)$ in Fig. 5a have been numerically fitted with the fit form $\rho(T)=A_{2}T^{2}$, and the fitted result of the low-temperature resistivity (black-dots) at $\delta=0.09$ is also plotted in Fig. 5a (inset), where the low-temperature resistivity is demonstrated clearly to be grown quadratically as the temperature is raised. For a better comparison, the corresponding experimental resultMirzaei13 of the low-temperature T-quadratic resistivity observed in the underdoped HgBa2CuO4+δ is also shown in Fig. 5b, where the most characteristic feature of the low-temperature resistivity in the underdoped regime is that it is perfectly quadratic down to the lowest achievable temperatures Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 . Apparently, (i) this characteristic feature of the T-quadratic behaviour of the low-temperature resistivity is the same in the theory and experiments Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 ; (ii) the magnitude of the low-temperature T-quadratic resistivity at a given doping and a given temperature is also qualitatively consistent with the corresponding experimental results in the underdoped regime Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , where different magnitudes at a given doping and a given temperature have been observed for different families of cuprate superconductors.
@@
Moreover, the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) $A_{2}$ grows as the doping is reduced. To see this doping dependence of the T-quadratic resistivity strength more clearly, we plot $A_{2}$ as a function of doping $\delta$ in Fig. 6a in comparison with the corresponding experimental resultsBarisic13 of the underdoped cuprate superconductors in Fig. 6b. It thus shows clearly that $A_{2}$ drops gradually as doping is enhanced in the underdoped regime, indicating that $A_{2}$ is roughly proportional to the inverse of the doping concentration. This tendency of the doping dependent $A_{2}$ is also in qualitative agreement with the corresponding experimental results of cuprate superconductors in the underdoped regime Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 . These results in Fig. 5 and Fig. 6 in the underdoped regime along with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap.
@@
The low-temperature resistivity is mainly determined by the umklapp scattering rate $\gamma(\theta,T)$ in Eq. (27), however, this umklapp scattering rate $\gamma(\theta,T)$ is in turn directly connected with the kernel function $F(\theta,\theta^{\prime})$ (then the probability weight of the umklapp scattering) in Eq. (28) as we have mentioned above. For the further understanding of the nature of the umklapp scattering between electrons by the exchange of the full effective spin propagator, we first analyse the exotic feature of the kernel function $F(\theta,\theta^{\prime})$.
Figure 7: (Color online) Surface plot of the kernel function $F(\theta,\theta^{\prime})/F(\theta_{\rm AN},\theta^{\prime}_{\rm AN})$ at $\delta=0.09$ with $T=0.05J$, where AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively, while $F(\theta_{\rm AN},\theta^{\prime}_{\rm AN})$ is the magnitude of $F(\theta,\theta^{\prime})$ at the antinode.
@@
Figure 8: Umklapp scattering rate $\gamma(T)/\Delta\gamma$ at the antinode as a function of temperature for $\delta=0.09$, where $\Delta\gamma=\gamma(T_{0})-\gamma(0)$, with $T_{0}=0.13J$. The inset shows the detail of the temperature dependence of $\gamma(T)$ at the antinode in the far lower temperature region.
@@
In Fig. 7, we plot $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where the probability weight of the electron umklapp scattering is very strongly anisotropic in momentum space. In particular, as in the case of the overdoped strange-metal phaseMa23 , the almost all of the probability weight of the electron umklapp scattering in the underdoped pseudogap phase is also concentrated at around the antinodal region, leading to that the strongest umklapp scattering occurs at around this antinodal region. However, the weakest umklapp scattering appears at around the tips of the Fermi arcs, since the probability weight of the electron umklapp scattering there is almost none. Moreover, the strength of the electron umklapp scattering at around the nodal region is much weaker than that at around the antinodal region, since a very small amount of the probability weight of the electron umklapp scattering is concentrated at around the nodal region. This special momentum dependent distribution of the probability weight of the electron umklapp scattering in Fig. 7 together with the recent discussions of the momentum-dependent distribution of the probability weight of the electron umklapp scattering in the overdoped regimeMa23 therefore confirm that in the whole doping regime, the dominant contribution to the scattering rate mainly arises from the umklapp scattering between the relatively slow electrons at around the antinodal region.
@@
We now turn to explore the unusual evolution of the umklapp scattering rate (27) with temperature in the underdoped pseudogap phase. We have made a series of calculations for the umklapp scattering rate $\gamma(T)$ at different Fermi angles, and the result of $\gamma(T)$ as a function of temperature for $\delta=0.09$ at the antinode is plotted in Fig. 8, where the inset shows the detail of the evolution of $\gamma(T)$ with temperature at the antinode in the far lower temperature region. Our numerical fit demonstrates that $\gamma(T)$ is purely T-quadratic in the low-temperature region, i.e., it increases quadratically with temperature as the temperature is increased. In particular, as the case of the resistivity shown in the inset of Fig. 5a, this $\gamma(T)$ approaches zero as the temperature approaches zero. Moreover, although the magnitude of $\gamma(\theta,T)$ is strongly anisotropic in momentum space, the low-temperature T-quadratic behaviour of $\gamma(\theta,T)$ appears at an any given Fermi angle $\theta$. Comparing this result in Fig. 8 with the corresponding results of the low-temperature T-quadratic resistivity in Fig. 5a, it thus shows that the low-temperature T-quadratic behaviour of $\gamma(T)$ together with the temperature region are the same as the corresponding behaviour and region in the resistivity $\rho(T)$, which therefore confirms that the low-temperature T-quadratic resistivity with the corresponding temperature region is mainly governed by the low-temperature T-quadratic umklapp scattering rate with the corresponding temperature region.
@@
Now we give an explanation to show why the low-temperature resistivity has a T-quadratic behaviour in the underdoped pseudogap phase, with a dramatic switch to the T-linear behaviour in the overdoped strange-metal phase? The nature of the umklapp scattering rate in Eq. (27) is mainly determined by the nature of the kernel function $F(\theta,\theta^{\prime})$ in Eq. (28), however, this kernel function $F(\theta,\theta^{\prime})$ in the underdoped pseudogap phase is proportional to the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (19). From the full spin propagator in Eq. (8), the full spin spectral function is derived straightforwardly as $A_{\rm spin}({\bf k},\omega)=A^{(1)}_{\rm spin}({\bf k},\omega)+A^{(2)}_{\rm spin}({\bf k},\omega)$, where the components $A^{(1)}_{\rm spin}({\bf k},\omega)$ and $A^{(2)}_{\rm spin}({\bf k},\omega)$ are given by,
@@
respectively. However, during the calculation, we have found that the spectral weight in the component $A^{(1)}_{\rm spin}({\bf k},\omega)$ is several orders of magnitude greater than the corresponding spectral weight in the component $A^{(2)}_{\rm spin}({\bf k},\omega)$, leading to that the electron umklapp scattering is mainly mediated by the spin excitations from the component of $A^{(1)}_{\rm spin}({\bf k},\omega)$.
On the other hand, the spin excitations with the full spin excitation dispersion $\bar{\omega}_{1{\bf k}}$ around the AF wave vector ${\bf k}_{\rm A}=[\pm\pi,\pm\pi]$ have the largest density of states. With these special properties of the spin excitations in the underdoped pseudogap phase, the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (24) can be reduced approximately as,
@@
with $b_{\bf q}=d^{2}\bar{\omega}_{1{\bf q}}/(d^{2}q)$. The above results in Eq. (35) show that as in the case of the the overdoped strange-metal phaseMa23 , the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (34) scales with $p^{2}$, and then when the electron umklapp scattering kicks in, the energy scale is proportional to $\Delta^{2}_{p}$ due to the presence of this $p^{2}$ scaling in the full effective spin propagator (34). In this case, $T_{\rm scale}=\bar{b}\Delta^{2}_{p}$ can be identified as the low-temperature scale, with the average value $\bar{b}=(1/N)\sum\limits_{{\bf q}\in\{{\bf k}_{\rm A}\}}b({\bf q})$ that is a constant at a given doping, where the summation ${\bf q}\in\{{\bf k}_{\rm A}\}$ is restricted to the extremely small area $\{{\bf k}_{\rm A}\}$ around the ${\bf k}_{\rm A}$ point of BZ. However, this low-temperature scale $T_{\rm scale}$ is strongly doping dependent. To see this doping dependence of $T_{\rm scale}$ more clearly, we plot $T_{\rm scale}$ as a function of doping in Fig. 9, where $T_{\rm scale}$ is relatively high at around the slightly underdoped region, then it decreases when the doping is increased in the heavily underdoped region, and is reduced to a very low temperature in the overdoped regime. Moreover, $T_{\rm scale}$ as a function of doping in Fig. 9 in the underdoped regime presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$ shown in Fig. 2b, suggesting a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase.
@@ IV Summary and discussion
Within the framework of the kinetic-energy-driven superconductivity, we have rederived the full spin propagator, the full charge-carrier propagator, and the full electron propagator in the normal-state of cuprate superconductors, where the spin self-energy (then the spin pseudogap) is obtained explicitly in terms of the collective charge-carrier mode in the particle-hole channel, and the charge-carrier normal self-energy (then the charge-carrier pseudogap) is obtained explicitly in terms of the spin excitation mode, while the electron normal self-energy (then the normal-state pseudogap) originated from the charge-carrier self-energy (then the charge-carrier pseudogap) is due to the charge-spin recombination. Moreover, we have also shown that (i) the spin excitation energy dispersion is anisotropically renormalized due to the momentum dependence of the spin pseudogap. In particular, the antinodal spin pseudogap effect is particularly notable in the slightly underdoped region, and then this antinodal spin pseudogap rapidly decreases with the increase of doping in the heavily underdoped region, eventually abrupt disappearing at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. By virtue of this full spin propagator and the reconstructed EFS, we have investigated the low-temperature electrical transport in the underdoped pseudogap phase of cuprate superconductors, where the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived within the framework of the Boltzmann transport theory. Our results show that the dominant contribution to the low-temperature resistivity mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap. Moreover, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime, which suggests a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$. In the underdoped regime, the resistivity exhibts a T-quadratic behaviour in the low-temperature region below $T_{\rm scale}$, where the strength of the T-quadratic resistivity decreases with the increase of doping. However, in the overdoped regime, the resistivity is T-linear in the low-temperature region above $T_{\rm scale}$. The current results together with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped regime therefore show that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependence of the spin pseudogap lowers the density of states of the spin excitations at around the antinodal region in the response to the intense umklapp scattering, which reduces the strength of the umklapp scattering from electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. Concomitantly, the low-temperature resistivity exhibits a dramatic switch from the T-linear behaviour in the overdoped strange-metal phase to the T-quadratic behaviour in the underdoped pseudogap phase.
It should be emphasized that the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (24) is obtained in the underdoped regime, where the AFLRO correlation is absent, although the AFSRO correlation survives from the underdoped regime to the overdoped regime Fujita12 as we have mentioned in subsection II.4. In this case, the umklapp scattering between electrons in Eq. (18) by the exchange of this full effective spin propagator is better suited for the description of the electrical transport in the underdoped regime. However, in the extremely light-doped regime ($\delta\leq 0.05$), the AFLRO correlation has been identified by INS, muon spin rotation, and other measurements Kastner98 ; Fujita12 ; Vaknin87 ; Brewer88 . Moreover, the strength of the AFLRO correlation as a function of the doping concentration has been established by NMR measurementsKitaoka88 , where AFLRO is destructed by a few percent of the doping concentration ($\sim 0.05$). In this extremely light-doped regime, the electronic state of the system with some special features appear, in particular, the low-temperature resistivity shows a upturn Ando01 ; Ando04a ; Ando04 . Although the mechanism causing this resistivity upturn remains unclear, it is possible that this striking resistivity upturn can be also described in terms of the umklapp scattering between electrons by the exchange of the effective spin propagator, where the spin propagator should give a suitable description of the special magnetic properties with the AFLRO correlation. These and the related issues are under investigation now.